# F5-TTS: A Fairytaler that Fakes Fluent and Faithful Speech with Flow Matching

### v2

@@ 4 Experimental Setup
Datasets We utilize the in-the-wild multilingual speech dataset Emilia [57] to train our base models. After simply filtering out transcription failure and misclassified language speech, we retain approximately 95K hours of English and Chinese data. We also trained small models for ablation study and architecture search on WenetSpeech4TTS [58] Premium subset, consisting of a 945 hours Mandarin corpus. Base model configurations are introduced below, and small model configurations are in Appendix B.1. Three test sets are adopted for evaluation, which are LibriSpeech-PC test-clean [59], Seed-TTS test-en [39] with 1088 samples from Common Voice [60], and Seed-TTS test-zh with 2020 samples from DiDiSpeech [61] [^3] https://github.com/BytedanceSpeech/seed-tts-eval. Most of the previous English-only models are evaluated on different subsets of LibriSpeech test-clean while the used prompt list is not released, which makes fair comparison difficult. Thus we build and release a 4-to-10-second LibriSpeech-PC subset with 1127 samples to facilitate community comparisons.
@@
Training Our base models are trained to 1.2M updates with a batch size of 307,200 audio frames (0.91 hours), for over one week on 8 NVIDIA A100 80G GPUs. The AdamW optimizer [62] is used with a peak learning rate of 7.5e-5, linearly warmed up for 20K updates, and linearly decays over the rest of the training. We set 1 for the max gradient norm clip. The F5-TTS base model has 22 layers, 16 attention heads, 1024/2048 embedding/feed-forward network (FFN) dimension for DiT; and 4 layers, 512/1024 embedding/FFN dimension for ConvNeXt V2; in total 335.8M parameters. The reproduced E2 TTS, a 333.2M flat U-Net equipped Transformer, has 24 layers, 16 attention heads, and 1024/4096 embedding/FFN dimension. Both models use RoPE as mentioned in Sec.3.2, a dropout rate of 0.1 for attention and FFN, the same convolutional position embedding as in Voicebox[14].
@@
We directly use alphabets and symbols for English, use jieba [^4] https://github.com/fxsjy/jieba and pypinyin [^5] https://github.com/mozillazg/python-pinyin to process raw Chinese characters to full pinyins. The character embedding vocabulary size is 2546, counting in the special filler token and all other language characters exist in the Emilia dataset as there are many code-switched sentences. For audio samples we use 100-dimensional log mel-filterbank features with 24 kHz sampling rate and hop length 256. A random 70% to 100% of mel frames is masked for infilling task training. For CFG (Sec.2.2) training, first the masked speech input is dropped with a rate of 0.3, then the masked speech again but with text input together is dropped with a rate of 0.2. We assume that the two-stage control of CFG training may have the model learn more with text alignment.
@@
Inference The inference process is mainly elaborated in Sec.3.1. We use the Exponential Moving Averaged (EMA) [63] weights for inference, and the Euler ODE solver for F5-TTS (midpoint for E2 TTS as described in [36]). We use the pretrained vocoder Vocos [51] to convert generated log mel spectrograms to audio signals.
@@
Baselines We compare our models with leading TTS systems including, (mainly) autoregressive models: VALL-E 2 [13], MELLE [29], FireRedTTS [64] and CosyVoice [65]; non-autoregressive models: Voicebox [14], NaturalSpeech 3 [12], DiTTo-TTS [35], MaskGCT [66], Seed-TTSDiT [39] and our reproduced E2 TTS [36]. Details of compared models see Appendix A.
@@
Metrics We measure the performances under cross-sentence task. The model is given a reference text, a short speech prompt, and its transcription, and made to synthesize a speech reading the reference text mimicking the speech prompt speaker. In specific, we report Word Error Rate (WER) and speaker Similarity between generated and the original target speeches (SIM-o) for objective evaluation. For WER, we employ Whisper-large-v3 [67] to transcribe English and Paraformer-zh [68] for Chinese, following [39]. For SIM-o, we use a WavLM-large-based [69] speaker verification model to extract speaker embeddings for calculating the cosine similarity of synthesized and ground truth speeches. We use Comparative Mean Opinion Scores (CMOS) and Similarity Mean Opinion Scores (SMOS) for subjective evaluation. For CMOS, human evaluators are given randomly ordered synthesized speech and ground truth, and are to decide how higher the naturalness of the better one surpasses the counterpart, w.r.t. prompt speech. For SMOS, human evaluators are to score the similarity between the synthesized and prompt.
@@ 5 Experimental Results
Tab.5 and 5 show the main results of objective and subjective evaluations. We report the average score of three random seed generation results with our model and open-sourced baselines. We use by default a CFG strength of 2 and a Sway Sampling coefficient of $-1$ for our F5-TTS.
Table 1: Results on LibriSpeech test-clean and LibriSpeech-PC test-clean. The boldface indicates the best result, and * denotes the score reported in baseline papers with different subsets for evaluation. The Real-Time Factor (RTF) is computed with the inference time of 10s speech. #Param. stands for the number of learnable parameters and #Data refers to the used training dataset in hours.
Model             #Param. #Data(hrs) WER(%)↓ SIM-o ↑ RTF ↓
@@
Ground Truth (2.2 hours subset) - - 2.2* 0.754* -
[…]
MELLE [29] - 50K EN 2.10* 0.625* 0.549*
MELLE-R2 [29] - 50K EN 2.14* 0.608* 0.276*
Voicebox [14] 330M 60K EN 1.9* 0.662* 0.64*
DiTTo-TTS [35] 740M 55K EN 2.56* 0.627* 0.162*
Ground Truth (40 samples subset) - - 1.94* 0.68* -
[…]
NaturalSpeech 3 [12] 500M 60K EN 1.94* 0.67* 0.296*
MaskGCT [66] 1048M 100K Multi. 2.634* 0.687* -
@@
Ground Truth (1127 samples 2 hrs) - - 2.23 0.69 -
Vocoder Resynthesized - - 2.32 0.66 -
[…]
FireRedTTS [64] $\sim$580M 248K Multi. 2.69 0.47 0.84
E2 TTS (32 NFE) [36]  [^3] M 100K Multi. 2.95 0.69 0.68
[…]
@@
For English zero-shot evaluation, the previous works are hard to compare directly as they use different subsets of LibriSpeech test-clean [70]. Although most of them claim to filter out 4-to-10-second utterances as the generation target, the corresponding prompt audios used are not released. Therefore, we build a 4-to-10-second sample test set based on LibriSpeech-PC [59] which is an extension of LibriSpeech with additional punctuation marks and casing. To facilitate future comparison, we release the 2-hour test set with 1,127 samples, sourced from 39 speakers (LibriSpeech-PC missing one speaker).
F5-TTS achieves a WER of 2.42 on LibriSpeech-PC test-clean with 32 NFE and Sway Sampling, demonstrating its robustness in zero-shot generation. Inference with 16 NFE, F5-TTS gains an RTF of 0.15 while still supporting high-quality generation with a WER of 2.53. It is clear that the Sway Sampling strategy greatly improves performance. The reproduced E2 TTS shows an excellent speaker similarity (SIM) but much worse WER in the zero-shot scenario, indicating the inherent deficiency of alignment robustness.
From the evaluation results on the Seed-TTS test sets, F5-TTS behaves similarly with a close WER to ground truth and comparable SIM scores. It produces smooth and fluent speech in zero-shot generation with a CMOS of 0.31 (0.21) and SMOS of 3.89 (3.83) on Seed-TTS test-en (test-zh), and surpasses some baseline models trained with larger scales. It is worth mentioning that Seed-TTS with the best result is trained with orders of larger model size and dataset (several million hours) than ours. As stated in Sec.3.1, we simply estimate duration based on the ratio of the audio prompt’s transcript length and the text prompt length. If providing ground truth duration, F5-TTS with 32 NFE and Sway Sampling will have WER of 1.74 for test-en and 1.53 for test-zh while maintaining the same SIM, indicating a high upper bound. A robustness test on ELLA-V [15] hard sentences is further included in Appendix B.5.
@@
[…]
FireRedTTS [64]     3.82 0.46 -1.46 2.94
MaskGCT [66]     2.623* 0.717* - -
Seed-TTSDiT[39]     1.733* 0.790* - -
E2 TTS (32 NFE) [36]     2.19 0.71   0.06 3.81
[…]
@@
[…]
FireRedTTS [64]     1.51 0.63 -0.49 3.28
MaskGCT [66]     2.273* 0.774* - -
Seed-TTSDiT[39]     1.178* 0.809* - -
E2 TTS (32 NFE) [36]     1.97 0.73 -0.04 3.44
[…]
@@ 5.1 Ablation of Model Architecture
To clarify our F5-TTS’s efficiency and stress the limitation of E2 TTS. We conduct in-depth ablation studies. We trained small models to 800K updates (each on 8 NVIDIA RTX 3090 GPUs for one week), all scaled to around 155M parameters, on the WenetSpeech4TTS Premium 945 hours Mandarin dataset with half the batch size and the same optimizer and scheduler as base models. Details of small model configurations see Appendix B.1.
We first experiment with pure adaLN DiT (F5-TTS--Conv2Text), which fails to learn alignment given simply padded character sequences. Based on the concept of refining the input text representation to better align with speech modality, and keep the simplicity of system design, we propose to add jointly learned structure to the input context. Specifically, we leverage ConvNeXt’s capabilities of capturing local connections, multi-scale features, and spatial invariance for the input text, which is our F5-TTS. And we ablate with adding the same branch for input speech, denoted F5-TTS$+$Conv2Audio. We further conduct experiments to figure out whether the long skip connection and the pre-refinement of input text are beneficial to the counterpart backbone, i.e. F5-TTS and E2 TTS, named F5-TTS$+$LongSkip and E2 TTS$+$Conv2Text respectively. We also tried with the Multi-Modal DiT (MMDiT) [37] a double-stream joint-attention structure for the TTS task which learned fast and collapsed fast, resulting in severe repeated utterance with wild timbre and prosody. We assume that the pure MMDiT structure is far too flexible for rigorous task e.g. TTS which needs more faithful generation following the prompt guidance.
@@
Figure 2: Ablation studies on model architecture. Seed-TTS test-zh evaluation results of 155M small models trained with WenetSpeech4TTS Premium a 945 hours Mandarin Corpus.
Fig.2 shows the overall trend of small models’ WER and SIM scores evaluated on Seed-TTS test-zh. Trained with only 945 hours of data, F5-TTS (32 NFE w/o SS) achieves a WER of 4.17 and a SIM of 0.54 at 800K updates, while E2 TTS is 9.63 and 0.53. F5-TTS$+$Conv2Audio trades much alignment robustness (+1.61 WER) with a slightly higher speaker similarity (+0.01 SIM), which is not ideal for scaling up. We found that the long skip connection structure can not simply fit into DiT to improve speaker similarity, while the ConvNeXt for input text refinement can not directly apply to the flat U-Net Transformer to improve WER as well, both showing significant degradation of performance. To further analyze the unsatisfactory results with E2 TTS, we studied the consistent failure (unable to solve with re-ranking) on a 7% of the test set (WER$>$50%) all along the training process. We found that E2 TTS typically struggles with around 140 samples which we speculate to have a large distribution gap with the train set, while F5-TTS easily tackles this issue.
We investigate the models’ behaviors with different input conditions to illustrate the advantages of F5-TTS further and disclose the possible reasons for E2 TTS’s deficiency. See from Tab.4 in Appendix B.2, providing the ground truth duration allows more gains on WER for F5-TTS than E2 TTS, showing its robustness in alignment. By dropping the audio prompt, and synthesizing speech solely with the text prompt, E2 TTS is free of failures. This phenomenon implied a deep entanglement of semantic and acoustic features within E2 TTS’s model design. From Tab.3 GFLOPs statistics, F5-TTS carries out faster training and inference than E2 TTS.
The aforementioned limitations of E2 TTS greatly hinder real-world application as the failed generation cannot be solved with re-ranking. Supervised fine-tuning facing out-of-domain data or a tremendous pretraining scale is mandatory for E2 TTS, which is inconvenient for industrial deployment. On the contrary, our F5-TTS better handles zero-shot generation, showing stronger robustness.
@@ 5.2 Ablation of Sway Sampling
It is clear from Fig.3 that a Sway Sampling with a negative $s$ improves the generation results. Further with a more negative $s$, models achieve lower WER and higher SIM scores. We additionally include comparing results on base models with and without Sway Sampling in Appendix B.4.
As stated at the end of Sec.3.2, Sway Sampling with $s<0$ scales more flow step toward early-stage inference ($t\to 0$), thus having CFM models capture more startup information to sketch the contours of target speech better. To be more concrete, we conduct a “leak and override" experiment. We first replace the Gaussian noise input $x_{0}$ at inference time with a ground-truth-information-leaked input $(1-t^{\prime})x_{0}+t^{\prime}x_{ref}^{\prime}$, where $t^{\prime}=0.1$ and $x_{ref}^{\prime}$ is a duplicate of the audio prompt mel features. Then, we provide a text prompt different from the duplicated audio transcript and let the model continue the subsequent inference (skip the flow steps before $t^{\prime}$). The model succeeds in overriding leaked utterances and producing speech following the text prompt if Sway Sampling is used, and fails without. Uniformly sampled flow steps will have the model producing speech dominated by leaked information, speaking the duplicated audio prompt’s context. Similarly, a leaked timbre can be overridden with another speaker’s utterance as an audio prompt, leveraging Sway Sampling.
@@
Figure 3: Probability density function of Sway Sampling on flow step $t$ with different coefficient $s$ (left), and small models’ performance on Seed-TTS test-zh with Sway Sampling (right).

### v3

@@ 4 Experimental Setup
Datasets We utilize the in-the-wild multilingual speech dataset Emilia (He et al., 2024) to train our base models. After simply filtering out transcription failure and misclassified language speech, we retain approximately 95K hours of English and Chinese data. We also trained small models for ablation study and architecture search on WenetSpeech4TTS (Ma et al., 2024) Premium subset, consisting of a 945 hours Mandarin corpus. Base model configurations are introduced below, and small model configurations are in Appendix B.1. Three test sets are adopted for evaluation, which are LibriSpeech-PC test-clean (Meister et al., 2023), Seed-TTS test-en (Anastassiou et al., 2024) with 1088 samples from Common Voice (Ardila et al., 2019), and Seed-TTS test-zh with 2020 samples from DiDiSpeech (Guo et al., 2021) [^2] https://github.com/BytedanceSpeech/seed-tts-eval. Most of the previous English-only models are evaluated on different subsets of LibriSpeech test-clean while the used prompt list is not released, which makes fair comparison difficult. Thus we build and release a 4-to-10-second LibriSpeech-PC subset with 1127 samples to facilitate community comparisons.
@@
Training Our base models are trained to 1.2M updates with a batch size of 307,200 audio frames (0.91 hours), for over one week on 8 NVIDIA A100 80G GPUs. The AdamW optimizer (Loshchilov, 2017) is used with a peak learning rate of 7.5e-5, linearly warmed up for 20K updates, and linearly decays over the rest of the training. We set 1 for the max gradient norm clip. The F5-TTS base model has 22 layers, 16 attention heads, 1024/2048 embedding/feed-forward network (FFN) dimension for DiT; and 4 layers, 512/1024 embedding/FFN dimension for ConvNeXt V2; in total 335.8M parameters. The reproduced E2 TTS, a 333.2M flat U-Net equipped Transformer, has 24 layers, 16 attention heads, and 1024/4096 embedding/FFN dimension. Both models use RoPE as mentioned in Sec.3.2, a dropout rate of 0.1 for attention and FFN, the same convolutional position embedding as in Voicebox(Le et al., 2024).
@@
We directly use alphabets and symbols for English, use jieba [^3] https://github.com/fxsjy/jieba and pypinyin [^4] https://github.com/mozillazg/python-pinyin to process raw Chinese characters to full pinyins. The character embedding vocabulary size is 2546, counting in the special filler token and all other language characters exist in the Emilia dataset as there are many code-switched sentences. For audio samples we use 100-dimensional log mel-filterbank features with 24 kHz sampling rate and hop length 256. A random 70% to 100% of mel frames is masked for infilling task training. For CFG (Sec.2.2) training, first the masked speech input is dropped with a rate of 0.3, then the masked speech again but with text input together is dropped with a rate of 0.2 (Le et al., 2024). We assume that the two-stage control of CFG training may have the model learn more with text alignment.
@@
Inference The inference process is mainly elaborated in Sec.3.1. We use the Exponential Moving Averaged (EMA) (Karras et al., 2024) weights for inference, and the Euler ODE solver for F5-TTS (midpoint for E2 TTS as described in Eskimez et al. (2024)). We use the pretrained vocoder Vocos (Siuzdak, 2023) to convert generated log mel spectrograms to audio signals.
@@
Baselines We compare our models with leading TTS systems including, (mainly) autoregressive models: VALL-E 2 (Chen et al., 2024), MELLE (Meng et al., 2024), FireRedTTS (Guo et al., 2024a) and CosyVoice (Du et al., 2024b); non-autoregressive models: Voicebox (Le et al., 2024), NaturalSpeech 3 (Ju et al., 2024), DiTTo-TTS (Lee et al., 2024), MaskGCT (Wang et al., 2024), Seed-TTSDiT (Anastassiou et al., 2024) and our reproduced E2 TTS (Eskimez et al., 2024). Details of compared models see Appendix A.
@@
Metrics We measure the performances under cross-sentence task (Wang et al., 2023a; Le et al., 2024). We report Word Error Rate (WER) and speaker Similarity between generated and the original target speeches (SIM-o (Le et al., 2024)) for objective evaluation. For WER, we employ Whisper-large-v3 (Radford et al., 2023) to transcribe English and Paraformer-zh (Gao et al., 2023b) for Chinese, following (Anastassiou et al., 2024). For SIM-o, we use a WavLM-large-based (Chen et al., 2022) speaker verification model to extract speaker embeddings for calculating the cosine similarity of synthesized and ground truth speeches. We use Comparative Mean Opinion Scores (CMOS) and Similarity Mean Opinion Scores (SMOS) for subjective evaluation. Details of subjective evaluations can be found in Appendix C.
@@ 5 Experimental Results
Tab.1 and 2 show the main results of objective and subjective evaluations. We report the average score of three random seed generation results with our model and open-sourced baselines. We use by default a CFG strength of 2 and a Sway Sampling coefficient of 1-1 for our F5-TTS.
Model        #Param.    #Data   WER(%)↓ SIM-o↑ RTF↓
@@
Ground Truth (2.2 hours subset) 2.2 0.754 -
VALL-E 2       - 50K EN 2.44 0.643 0.732
MELLE        - 50K EN 2.10 0.625 0.549
MELLE-R2       - 50K EN 2.14 0.608 0.276
Voicebox       330M 60K EN 1.9 0.662 0.64
DiTTo-TTS      740M 55K EN 2.56 0.627 0.162
Ground Truth (40 samples subset) 1.94 0.68 -
Voicebox       330M 60K EN 2.03 0.64 0.64
NaturalSpeech 3    500M 60K EN 1.94 0.67 0.296
MaskGCT      1048M 100K Multi. 2.634 0.687 -
@@
Ground Truth (1127 samples 2 hrs) 2.23 0.69 -
Vocoder Resynthesized 2.32 0.66 -
CosyVoice      $\sim$300M 170K Multi. 3.59 0.66 0.92
FireRedTTS     $\sim$580M 248K Multi. 2.69 0.47 0.84
E2 TTS (32 NFE)    [^3] M 100K Multi. 2.95 0.69 0.68
F5-TTS (16 NFE)   336M 100K Multi. 2.53 0.66 0.15
@@
Table 1: Comparison results on LibriSpeech(-PC) test-clean. The Real-Time Factor (RTF) is computed with the inference time of 10s speech on NVIDIA RTX 3090. #Param. stands for the number of learnable parameters and #Data refers to the used training dataset in hours.
F5-TTS achieves a WER of 2.42 on LibriSpeech-PC test-clean with 32 NFE, demonstrating its robustness in zero-shot generation. Inference with 16 NFE, F5-TTS gains an RTF of 0.15 while still supporting high-quality generation with a WER of 2.53. The reproduced E2 TTS shows an excellent speaker similarity (SIM) but much worse WER in the zero-shot scenario, indicating the inherent deficiency of alignment robustness.
From the evaluation results on the Seed-TTS test sets, F5-TTS behaves similarly with a close WER to ground truth and comparable SIM scores. It produces smooth and fluent speech in zero-shot generation with a CMOS of 0.31 (0.21) and SMOS of 3.89 (3.83) on Seed-TTS test-en (test-zh), and surpasses some baseline models trained with larger scales. As stated in Sec.3.1, we simply estimate duration based on the ratio of the audio prompt’s transcript length and the text prompt length, rather than relying on an extra duration predictor. It is also worth mentioning that Seed-TTS with the best result is trained with orders of larger model size and dataset (several million hours) than ours.
A robustness test on ELLA-V (Song et al., 2024) hard sentences is further included in Appendix B.5. The ablation of vocoders and additional evaluation with a non-PC test set are in Appendix B.6. An analysis of training stability with varying data scales is in Appendix B.7.
@@
CosyVoice    3.39 0.64   0.02 3.64
FireRedTTS    3.82 0.46 -1.46 2.94
MaskGCT    2.623* 0.717* - -
Seed-TTSDiT    1.733* 0.790* - -
E2 TTS (32 NFE)    2.19 0.71   0.06 3.81
F5-TTS (16 NFE)    1.89 0.67   0.16 3.79
@@
CosyVoice    3.10 0.75 -0.06 3.54
FireRedTTS    1.51 0.63 -0.49 3.28
MaskGCT    2.273* 0.774* - -
Seed-TTSDiT    1.178* 0.809* - -
E2 TTS (32 NFE)    1.97 0.73 -0.04 3.44
F5-TTS (16 NFE)    1.74 0.75   0.02 3.72
@@ 5.1 Ablation of Model Architecture
To clarify our F5-TTS’s efficiency and stress the limitation of E2 TTS. We conduct in-depth ablation studies. We trained small models (all around 155M parameters) to 800K updates on the WenetSpeech4TTS Premium 945 hours Mandarin dataset with half the batch size compared to base models. Configuration details see Appendix B.1.
@@
Figure 2: Seed-TTS test-zh evaluation results of 155M small models trained with WenetSpeech4TTS Premium a 945 hours Mandarin Corpus.
Fig.2 shows the overall trend of productive small models’ WER and SIM scores evaluated on Seed-TTS test-zh. F5-TTS (32 NFE w/o SS) achieves a WER of 4.17 and a SIM of 0.54 at 800K updates, while E2 TTS is 9.63 and 0.53. We were first faced with unsatisfactory results with reproduced E2 TTS in Mandarin. On the one hand, E2 TTS converges slowly; on the other hand, we found that E2 TTS consistently failed (unable to solve with re-ranking) on 7% test samples (WER>>50%) all along the training process, speculated of larger distribution gap with train set. To disclose the possible reasons for E2 TTS’s deficiency, we investigate the models’ behaviors with different inputs. See from Tab.4 in Appendix B.2, by dropping the audio prompt, and synthesizing speech solely with the text prompt, E2 TTS is free of failures (F5-TTS also benefits but because of more standard output that is easier to be recognized by the ASR model). This phenomenon implied a deep entanglement of semantic and acoustic features within E2 TTS’s model design, and it greatly hinders real-world application as the failed generation cannot be solved with re-ranking. Supervised fine-tuning facing out-of-domain data or a tremendous pretraining scale under the slow convergence speed is mandatory for E2 TTS, which is inconvenient for industrial deployment and a crushing burden for individuals.
From Tab.3 GFLOPs, structures without many skip connections natively allow faster training and inference. However, pure adaLN DiT (F5-TTS-Conv2Text) failed to learn alignment given simply padded character sequences, while MMDiT (Esser et al., 2024) learned fast and collapsed fast, resulting in severe repeated utterance with wild timbre and prosody. We assume that the pure MMDiT structure is far too flexible for TTS task that requires faithful generation following guidance. Thus we focus on enhancing the modeling ability of DiT. Based on the concept of refining the input text representation to better align with speech modality, and keep the simplicity of system design, F5-TTS is proven effective. F5-TTS easily handles zero-shot generation, showing stronger robustness.
We further ablate with adding the same branch for input speech (F5-TTS++Conv2Audio), and also conduct experiments to figure out whether the long skip connection and the refinement of input text are beneficial to the counterpart backbone, i.e. F5-TTS and E2 TTS, named F5-TTS++LongSkip and E2 TTS++Conv2Text respectively. From Fig.2, F5-TTS++Conv2Audio trades much alignment robustness (+1.61 WER) with a slightly higher speaker similarity (+0.01 SIM). The long skip connection structure can not simply fit into DiT to improve speaker similarity, while the ConvNeXt for input text refinement can not directly apply to the flat U-Net Transformer to improve WER as well, both showing significant degradation of performance.
@@ 5.2 Ablation of Sway Sampling
It is clear from Fig.3 that a Sway Sampling with more negative $s$ value further improves performance. Appendix B.3 with massive ablation results on base models, provides more evidence of the effectiveness of the proposed strategy.
Figure 3: The probability density function of Sway Sampling with different coefficient $s$, and small models’ corresponding performance on Seed-TTS test-zh.
To be more concrete and intuitive, we conduct a "leak and override" experiment. We first replace the Gaussian noise input $x_{0}$ at inference time with a ground-truth-information-leaked input $(1-t^{\prime})x_{0}+t^{\prime}x_{ref}^{\prime}$, where $t^{\prime}=0.1$ and $x_{ref}^{\prime}$ is a duplicate of the audio prompt mel features. Then, we provide a text prompt different from the duplicated audio transcript and let the model continue the subsequent inference (skip the flow steps before $t^{\prime}$). The model succeeds in overriding leaked utterances and producing speech following the text prompt if Sway Sampling is used, and fails without. Uniformly sampled flow steps will have the model producing speech dominated by leaked information, speaking the duplicated audio prompt’s context. Similarly, a leaked timbre can be overridden with another speaker’s utterance as an audio prompt, leveraging Sway Sampling.